Problem : 7+w - 10 = 2w
Question : What is w ?

Found 2 solutions by sophxmai, mananth:
Answer by sophxmai(62)

7+w - 10 = 2w
7-10=2w-w
-3=w

Answer by mananth(16946)

7+w - 10 = 2w
7-10+w=2w
-3+w=2w
add 3 to both sides
-3+3+w=2w+3
w=2w+3
add -2w to both sides
w-2w=-2w+2w+3
-w=+3
/-1
w=-3
...
m.ananth@hotmail.ca